Restart Night #1 UT 2018-09-02

Observers: J. Hill, B. Rothberg (Tucson)

OSA: S. Allanson, assisted by J. Wiliams

ISp: E. Solheid

Mtn Manager: P. Hartley

Before Opening

LBC functional checks completed (rotator and filter wheels).

LBC test biases and dome flats (not for use) completed.

LUCI functional testing continued.

After Sunset

LUCI functional testing continued.

Power bump at 09:57 (Local Time) - internet connectivity lost. LBC and LUCI testing halted.

Night Officially called at UT 07hr - telescope slewing to horizon due to onslaught of rain.

Article text.

-- %USERSIG{BarryRothberg - 2018-09-02}%

Comments


 
Topic revision: r1 - 02 Sep 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback